OpenStudy (anonymous):

Verify: 1.(cscx/cosx)-tan=cotx Solve: 2. cos^2x+sin(2x)-sinx-1=-sin^2x

OpenStudy (aravindg):

simplify csc x/cot x first

OpenStudy (anonymous):

1) left side: \[\frac{ \frac{ 1 }{sinx } }{ cosx }-\frac{ sinx }{ cosx}\] \[\frac{ ccosx-\sin^2cosx }{ \cos^2xsinx }\] \[\frac{ 1-\sin^2x }{ sinx cosx }\] \[\frac{ \cos^2x }{ sinx cosx}=\frac{ cosx }{sinx }=cotx\]

OpenStudy (anonymous):

@ASAAD123 how did you get from step 1 to step 2 ?

OpenStudy (anonymous):

ya were did "ccosx-sin^2cosx" come from?

OpenStudy (anonymous):